I need help with how to solve these simultaneous equations. Could someone please show me the steps to get to the answer.

Sum of forces in the x direction is zero i.e. FCD COs 33.69-FAD COS 33.69-FDE COS 33.690-0 54.08xcos 33.69 - FAD coS 33.69 F cos 33.69... .(5) Sum of forces in the y direction is zero i.e FRD +Fcn sin 33.69 + FAD sin 33.69 - FDE sin 33.690 -20-54.08x sin 33.69 + FD sin 33.69-Fg sin 33.690 (6) By solving the equations (5) and (6), we get FAD 18.03 kN and FDE72 kN
Answer #1

we people generally use scientific calculator for calculation of equation.

-54.08xCos33.69o-FADCos33.69o-FDECos33.69o =0 ------(5)

-20-54.08xSin33.69o+FADSin33.69o-FDESin33.69o =0 --------(6)

divide whole term in equation (5) by Cos33.69o

-54.08-FAD-FDE =0

or,FAD+FDE = -54.08 -----(5A)

divide whole term in equation (6) by Sin33.69o

-20/(Sin33.69o)-54.08+FAD-FDE = 0  

Sin33.69o = 0.5547 ( you have to use scientific calculator for this,there is no other way)

-20/0.5547-54.08+FAD-FDE = 0

FAD-FDE = 90.1355 ------(6A)

now you can use scientific calculator for solving (5A) and (6A) or,

add (5A) and (6A) FAD+FDE +FAD-FDE= -54.08+ 90.1355

2FAD =36.0555

  FAD = 18.02775KN = 18.03KN

now put this value of FAD in equation (5A) or (6A),let's put in (5A)

18.03+FDE = -54.08

FDE = -72.11KN